# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

### v3

@@ Abstract
$\mathbf{h}$ $=\operatorname{Dropout}\left(\operatorname{ReLU}\left(\text{\bf W}_{1}\bm{e}\right)\right)$ (4)
$\mathbf{z}$ $=\operatorname{Dropout}\left(\operatorname{ReLU}\left(\text{\bf W}_{2}\bm{h}\right)\right)$ (5)
@@
Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $20\%$ for the linear MLP and $15\%$ for non-linear MLP.
@@
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. To conduct a more accurate and reliable evaluation, we use the original C-STS train set and re-annotated validation set. A 70-30 split is used for the re-annotated validation set, with randomly selected 70% of the data allocated for validation and the remaining 30% for testing. We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Scikit-learn 1.3.0 is used for PCA and ICA. Pytorch 2.0.1 with cuda 11.7 is used for MLP proejction. These settings are fixed across all experiments. For reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a linear MLP for CASE takes approximately 1.5 minutes, whereas a non-linear MLP requires about 5 minutes (wall-clock time).
@@
We evaluate the performance of sentence embedding models on two metrics: Spearman Rank correlation and Accuracy.
Spearman Rank Correlation:
We compute Spearman’s rank correlation between the similarity scores by CASE and the re-annotated human ratings on the test set.
Accuracy
Spearman correlation is highly sensitive to small variations in similarity scores, which can be affected by the noise in human annotations. Given the subjectivity and ambiguity of human ratings, we introduce Accuracy as a more robust alternative to assess whether the model correctly captures the relative impact of conditions on semantic similarity. Different from Spearman correlation, which considers actual similarity values, accuracy only evaluates whether the predicted similarity under the higher-rated condition $c_{\rm high}$ is greater than that under the lower-rated condition $c_{\rm low}$.
For each sentence pair $\left(s_{1},s_{2}\right)$, there exist two conditions $c_{1}$ and $c_{2}$ with corresponding human labelled similarity scores $y_{1}$ and $y_{2}$, where $y_{1}>y_{2}$.
Cosine similarity is computed between CASE under the same condition for the similarity score
$\operatorname{sim}_{c_{1}}=\operatorname{cos}\left({\rm CASE}(s_{1},c_{1}),{\rm CASE}(s_{2},c_{1})\right)$ (6)
$\operatorname{sim}_{c_{2}}=\operatorname{cos}\left({\rm CASE}(s_{1},c_{2}),{\rm CASE}(s_{2},c_{2})\right)$ (7)
A prediction is considered correct if
$\left(\operatorname{sim}_{c_{1}}-\operatorname{sim}_{c_{2}}\right)\left(y_{1}-y_{2}\right)>0,$ (8)
which evaluates whether the model’s predicted similarity ranking aligns with the human annotations. Then, the accuracy is given by (9)
$\frac{\sum_{i=1}^{N}1\left[\left(\operatorname{sim}_{c_{1}}^{(i)}-\operatorname{sim}_{c_{2}}^{(i)}\right)\left(y_{1}^{(i)}-y_{2}^{(i)}\right)>0\right]}{N},$ (9)
where $N$ is the total number of test instances, and $1[\cdot]$ is the indicator function.
@@
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM-based models, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM-based models, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we evaluate the effect of subtracting the condition embedding, $\bm{c}=f(c;I(\emptyset))$.
The test performance for different settings and models are shown in Table 1. For LLM-based embeddings, cond consistently reports higher Spearman correlation than sent, suggesting that embedding the condition given the sentence is more effective for C-STS measurement than embedding the sentence given the condition. Moreover, we see that subtracting $\bm{c}$ further improves performance both in terms of Spearman correlation and accuracy across all settings, except for accuracy in SFR. The former approach reduces the noise due to the tokens in a sentence, which are irrelevant to the given condition. For MLM-based embeddings, subtracting $\bm{c}$ also improves performance. It reduces condition-specific information that are not altered by the sentence, thus allowing CASE to focus on the information that varies between the two sentences being compared. Moreover, we discovered that subtracting $\bm{c}$ in a post-processing step improves isotropy of the embeddings as shown in Appendix B. This is in-line with prior work reporting a correlation between isotropy and improved performance in embedding models Rajaee and Pilehvar (2022); Su et al. (2021). The effect of pooling method on performance is discussed in Appendix C, where we find the latent pooling in NV to perform best. Therefore, we use the ($\operatorname{cond}-\bm{c}$) setting (which reports the best performance) for the six sentence encoders to conduct the subsequent experiments.
Model sent/cond? Spear. Acc.
NV sent - c 16.98 52.24
sent 22.07 59.89
cond - c 31.32 64.91
cond 27.02 48.02
SFR sent - c 19.54 57.52
sent 11.89 43.01
cond - c 20.38 52.51
cond 18.32 46.44
GTE sent - c 7.16 42.48
sent 7.16 37.20
cond - c 20.40 54.08
cond 16.58 45.12
E5 sent - c 11.08 42.21
sent 3.77 33.77
cond - c 12.01 42.74
cond 6.18 37.47
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59 37.46
$\operatorname{CONC}(c+s)$ 4.00 35.09
$\operatorname{CONC}(s+c)$ c 8.32 34.56
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ c 8.58 48.02
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37 39.81
$\operatorname{CONC}(c+s)$ 1.25 34.82
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ c 7.05 42.74
$\operatorname{CONC}(c+s)$ c 6.00 36.67
Table 1: Spearman and accuracy scores for different sentence embedding models and encoding settings.
(a) Test Spearman for linear MLP
(b) Test Spearman for non-linear MLP
@@
Figure 3: Spearman correlation on test set for different models over training steps with dimensionality 512. The $y$-axes of both subfigures are aligned, facilitating a direct comparison of the Spearman correlation coefficients across the two line charts, with the same colour for the same model. Best viewed in colour.
Figure 4: Spearman correlation coefficients of the four dimensionality reduction methods on the test set for NV-Embed-v2 (NV) over different dimensionalities.
Model Non-linear MLP Linear MLP PCA ICA
NV 69.30 69.95 32.04 19.13
SFR 62.85 59.22 19.86 -1.99
GTE 64.16 56.10 22.82 7.26
E5 62.12 47.03 11.43 -0.17
SimCSE_large 56.67 45.96 8.22 -4.89
SimCSE_base 56.60 39.54 6.19 -1.66
@@
Table 2: Spearman correlation of embedding models based on supervised and unsupervised dimensionality reduction methods with reduced dimensionality $512$.
@@
cos(s1, s2)   $0.5006\rightarrow 0.9016$
Condition 1: The color of the dress.  Rating: 1 Condition 2: The name of the game.  Rating: 5
cos(s1, s2; c1)   $0.4757\rightarrow 0.2522$ cos(s1, s2; c2)   $0.6233\rightarrow 0.9660$
s1: Two snow skiers with ski poles and snow skis, standing on top of a snow covered mountain with other skiers around them.
@@
cos(s1, s2)   $0.6318\rightarrow 0.7702$
Condition 1: The number of person.  Rating: 1 Condition 2: The type of job.        Rating: 5
cos(s1, s2; c1)    $0.6358\rightarrow 0.2788$ cos(s1, s2; c2)$0.7502\rightarrow 0.9539$
@@
cos(s1, s2)   $0.3988\rightarrow 0.5350$
Condition 1: The type of hobby.   Rating: 1 Condition 2: The type of location.      Rating: 5
cos(s1, s2; c1)   $0.4386\rightarrow 0.4886$ cos(s1, s2; c2)$0.4825\rightarrow 0.8582$
Table 3: Cosine similarity scores ($\cos$) for two sentences under different conditions are computed using NV embeddings (shown on the left side of the arrows) vs. using MLP projected $512$ embeddings (i.e. CASE embeddings given by (1)) (shown on the right side of the arrows). Compared to the unconditional similarity between the two sentences, conditional similarity scores computed using NV embeddings align well with the human ratings. Moreover, the conditional similarities are further appropriately amplified by the supervised projection. Further qualitative evaluations shown in Appendix E.
We show the training curves for our supervised MLPs in Figure 3. Overall, the non-linear MLPs achieve significantly higher Spearman correlation than the linear MLPs, except for NV, where the linear MLP performs slightly better than the non-linear MLP. Moreover, non-linear MLPs converge faster, typically reaching their peak performance within $20$ epochs, after which the performance declines due to overfitting. The performance of the linear MLPs, gradually increases and eventually converges as the training progresses. NV consistently performs the best for both linear and non-linear MLPs.
To explore the relationship between test performance and dimensionality, we use NV to evaluate our supervised dimensionality reduction methods (MLPs) as well as the unsupervised methods (PCA and ICA), as shown in Figure 4. Overall, the supervised MLPs achieve much higher Spearman correlations while significantly reducing the dimensionality. Their performance stabilises at relatively high levels as the dimensionality exceeds $200$ and peaks around $500$. In this way, we achieve an 8×8$\times$8 × compression of the sentence embeddings (original dimensionality $4096$) while maintaining high performance. In contrast, unsupervised PCA and ICA do not effectively capture the conditional semantics in sentence embeddings. While PCA preserves the performance of the original sentence embeddings, it does not improve the performance on the C-STS task. ICA, on the other hand, performs even worse, showing a consistent decline in performance as dimensionality increases, which could be attributable to the lack of sufficiently independent components in sentence embeddings.
We compare sentence encoders in Table 2 for a fixed (i.e. 512) dimensional projection from their original embeddings. NV obtains the highest Spearman coefficient across all sentence encoders and dimensionality reduction methods. GTE and SFR also achieve high Spearman coefficients under the non-linear MLPs. SimCSE_large and SimCSE_base have the lowest Spearman coefficients, regardless of the dimensionality reduction method being used. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task. Note that direct comparison between our method and other C-STS approaches is challenging due to differences in the train/test sets used in prior work. Nevertheless, we present our results alongside previous methods in Appendix D for completeness, which shows its superior performance.
@@
Similarity scores computed for three sentence pairs are shown in Table 3 for the original high-dimensional LLM-based embeddings and low-dimensional projected CASE. By comparing the similarity scores with and without the supervised projection, we can evaluate the ability of CASE to focus on the condition-related information. Compared to the unconditional similarity between the two sentences, conditional similarities scores computed using NV embeddings align well with the human ratings. For example, in the top row in Table 3, we see that the unconditional similarity between the two sentences reduces from $0.5006$ to $0.4747$ under $c_{1}$, while increasing to $0.6233$ under $c_{2}$. Moreover, the conditional similarities are further appropriately amplified by CASE using the supervised projection (i.e. decreasing to $0.2522$ under $c_{1}$, while increasing to $0.9660$ under $c_{2}$). Specifically, CASE reduces the similarity between the two sentences under the lower-rated condition, while increasing the same under the high-rated condition. This demonstrates that CASE is able to effectively capture the shift in conditional meaning between sentences under different conditions.
@@
We propose CASE, which encodes the condition under sentence by subtracting the condition in a post-processing step when measuring C-STS. Our findings show that LLM-based sentence embeddings consistently outperform MLM-based embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to reduce the dimensionality of LLM-based embeddings, while improving the performance in C-STS.

### v4

@@ Abstract
$\mathbf{z^{\prime}}$ $=\operatorname{Dropout}\left(\operatorname{LeakyReLU}\left(\text{\bf W}^{\prime}\bm{e}\right)\right)$ (4)
@@
A dropout layer is applied to reduce any overfitting Hinton et al. (2012). Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $15\%$ for non-linear FFN. Appendix B provides the ablation study for the architectures of FFNs.
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. Zhang et al. (2025) further identified the issues on both condition statements and ratings. They improved the dataset by first revising the condition statements and then re-annotating the similarity ratings for both the training and validation sets. To conduct a more accurate and reliable evaluation, we use the most recent revised and cleaned datasets by Zhang et al. (2025). [^2] https://huggingface.co/datasets/LivNLP/C-STS-Reannotated We split their validation set 70% vs. 30% randomly to a validation set of 1983 instances and a test set of 851 instances. In this C-STS dataset, there are 14176 unique instances with 4383 unique conditions. As for each sentence pair ($s_{1}$, $s_{2}$), there are two conditions ($c_{\rm high}$ and $c_{\rm low}$). There are 7088 unique sentence pairs, of which 5719 pairs have unique condition pairs ($c_{\rm high}$, $c_{\rm low}$). In other words, 19.3% of the sentence pairs share their conditions with other sentence pairs.
@@
We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Pytorch 2.0.1 with cuda 11.7 is used for FFN projection. These settings are fixed across all experiments. For reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a non-linear FFN for CASE takes less than 1 minute (wall-clock time).
@@
We evaluate the performance of a sentence encoder by the Spearman correlation coefficient between the cosine similarity scores computed using the embeddings produced by an encoder and the human similarity ratings on the test set.
@@
Model $I_{\text{iso}}$ (-c) $I_{\text{iso}}$
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 2: $I_{\text{iso}}$ values for each model. The left column of $I_{\text{iso}}$ (-c) lists values with the post-processing step of subtracting the embedding of condition c. $I_{\text{iso}}$ values close to 1 indicate high isotropy.
@@
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM embeddings, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM embeddings, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we evaluate the effect of subtracting the condition embedding, $\bm{c}=f(c;I(\emptyset))$.
(a) Test Spearman for linear FFN
(b) Test Spearman for non-linear FFN
@@
Figure 3: Spearman correlation on test set for different models over training steps with dimensionality 512. The $y$-axes of both subfigures are aligned, facilitating a direct comparison of the Spearman coefficients across the two line charts, with the same colour for the same model. Best viewed in colour.
The test performance for different settings and models are shown in Table 1. For LLM embeddings, cond consistently reports higher Spearman correlation than sent, suggesting that embedding the condition given the sentence is more effective for C-STS measurement than embedding the sentence given the condition. Moreover, we see that subtracting $\bm{c}$ further improves performance across all settings. The former approach reduces the noise due to the tokens in a sentence, which are irrelevant to the given condition. For MLM-based embeddings, subtracting $\bm{c}$ also improves performance. This operation removes the sentence-independent components of the condition, thereby isolating the contrastive semantics — the aspects of the condition that are actually altered by the sentence. As a result, CASE focuses on the semantic variation between the two sentences under the given condition, rather than the redundant information shared across all condition phrases.
Additionally, we discovered that subtracting $\bm{c}$ in a post-processing step improves isotropy of the embeddings (full details in Appendix C). We use the approximated isotropy metric $I_{\text{iso}}$ Durdy et al. (2023) as a numerically stable method to measure isotropy. $I_{\text{iso}}$ is estimated by randomly sampling unit vectors on the hypersphere and computing the ratio of their minimum to maximum alignment with the embeddings across random directions. Table 2 shows that the post-processing step of subtracting the condition $\bm{c}$ gives higher $I_{\text{iso}}$ values, indicating the improvement of isotropy. This makes subtle semantic differences between sentences under a given condition more distinguishable. This is in-line with prior work reporting a positive correlation between isotropy and improved performance in embedding models Rajaee and Pilehvar (2022); Su et al. (2021). The effect of pooling method on performance is discussed in Appendix D, where we find the latent pooling in NV to perform best. Therefore, we use the ($\operatorname{cond}-\bm{c}$) setting (which reports the best performance) for the six sentence encoders to conduct the subsequent experiments.
Model Non-linear FFN Linear FFN
NV 74.94 73.55
SFR 68.36 59.02
GTE 64.16 55.88
E5 65.07 48.25
SimCSE_large 63.64 48.15
SimCSE_base 61.52 42.22
@@
Table 3: Spearman correlation of embedding models with supervised methods at reduced dimensionality $512$.
Figure 4: Spearman correlation coefficients of our CASE for NV embeddings on the test set over different dimensionalities. The horizontal dashed line represents the performance of fine-tuned SimCSE-base (768-dimensional).
We show the training curves for our supervised FFNs in Figure 3. Overall, the non-linear FFNs achieve significantly higher Spearman correlation than the linear FFNs. Moreover, non-linear FFNs converge faster, typically reaching their peak performance within 20 epochs. The performance of the linear FFNs, gradually increases and eventually converges as the training progresses. NV consistently performs the best for both linear and non-linear FFNs.
We compare sentence encoders in Table 3 for a fixed (i.e. 512) dimensional projection from their original embeddings. NV obtains the highest Spearman coefficient across all sentence encoders and supervised methods. GTE and SFR also achieve high Spearman coefficients under the non-linear FFNs. SimCSE_large and SimCSE_base have the lowest Spearman coefficients. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task. Importantly, all embeddings are projected to the same lower dimensionality.
Method Spearman dim #params
Deshpande et al.
w/ SimCSE-large 79.51 1024 354.3M
w/ SimCSE-base 72.12 768 124.1M
CASE
w/ non-linear FFN 74.94 512 2.1M
Table 4: Comparison of Spearman correlation, embedding dimensionality, and trainable parameters of previously proposed methods and our CASE. #params denotes the number of trainable parameters.
@@
[…]
s1: Two snow skiers with ski poles and snow skis, standing on top of a mountain with other skiers around them.
@@
[…]
@@
[…]
Table 5: Example of similarity scores for two conditions applied to the same sentence pair, based on non-linear FFN with dimensionality $512$ on NV. The table shows how supervised FFN improves the CASE for C-STS task. cos(,)($\cdot$$\cdot$) denotes cosine similarity. Answer 1 and Answer 2 refer to the corresponding answers for Sentence 1 and Sentence 2 under conditions. The predicted similarity scores before and after applying supervised FFN are listed on the left and right of the arrow. That is, the similarity score for original high-dimensional LLM-based embeddings is on the left of the arrow, while the similarity score for CASE is on the right.
To study how the performance of the proposed CASE varies with the dimensionality of the supervised projection, we plot the Spearman correlation measured on the test split of the C-STS dataset in Figure 4 using NV embeddings. Because we are specifically interested in examining how CASE performs in the low-dimensional regime, which is attractive from a computational viewpoint, we consider dimensionalities in range [50, 1000]. Moreover, as a reference, we plot the performance of the previously proposed bi-encoder model by Deshpande et al. (2023) for the SimCSE-base model, which is the smallest dimensional (768) encoder used in their work. From Figure 4 we see that for dimensionalities greater than 200, CASE consistently outperforms SimCSE-base bi-encoder. This result has practical implications because it shows that we can reduce the high dimensionality of LLMs-based (NV) embeddings from 4096 to 512 with an 8X compression, while preserving its high performance. This result is consistent with Matryoshka Representation Learning (MRL) Kusupati et al. (2022), where they show that small-dimensional embeddings can maintain high performance in downstream tasks. MRL is applicable with our MLM and LLM embeddings to further reduce the dimensionality, while our FFNs project embeddings into a specific lower dimensionality rather than the nested granularities of MRL. Our results show that embeddings can be efficiently compressed in the bi-encoder setting on the C-STS task.
We compare CASE against the best bi-encoder models with fine-tuned SimCSE by Deshpande et al. (2023). As shown in Table 4, our proposed CASE with non-linear FFN achieves 74.9% Spearman correlation, outperforming fine-tuned SimCSE_base while using only 2.1M trainable parameters and a 512-dimensional embedding space. Compared with SimCSE_large, CASE maintains 94% of its performance, but requires 50% fewer dimensions and only 0.6% of its trainable parameters. This highlights that CASE is a simple but efficient method.
§ 4.3 Case Study
@@
To further illustrate how CASE captures condition-dependent semantics, Table 5 presents qualitative examples. Similarity scores are computed and compared for three sentence pairs for the original high-dimensional LLM-based embeddings and low-dimensional trained CASE. By comparing the similarity scores with and without the supervised projection, we can evaluate the ability of CASE to focus on the condition-related information. Compared to the unconditional similarity between the two sentences, conditional similarities scores computed using NV embeddings align well with the human ratings. For example, in the top row in Table 5, we see that the unconditional similarity between the two sentences reduces from $0.5006$ to $0.4747$ under $c_{1}$, while increasing to $0.6233$ under $c_{2}$. Moreover, the conditional similarities are further appropriately amplified by CASE using the supervised projection (i.e. decreasing to $0.2522$ under $c_{1}$, while increasing to $0.9660$ under $c_{2}$). Specifically, CASE reduces the similarity between the two sentences under the lower-rated condition, while increasing the same under the high-rated condition.
We also treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that CASE can effectively capture the shift in conditional meaning between sentences under different conditions.
Method Spearman Dim
Qwen3-8B (frozen)
     w/o FFN 24.59 4096
     w/ FFN (CASE) 71.64 512
Qwen3-8B (LoRA fine-tuned)
     w/o FFN 79.91 4096
     w/ FFN 79.11 512
Table 6: Comparison of fine-tuning. We report the Spearman correlation of the original and LoRA fine-tuned Qwen3-8B encoders, both with and without our supervised non-linear FFN.
§ 4.4 Comparision with Fine-tuning LLM Encoders
We investigate the effects of fine-tuning LLM encoders using LoRA Hu et al. (2022), and compare its performance with our CASE method. Specifically, we use the high-performing encoder Qwen3-Embedding-8B (Qwen3-8B) [^3] https://huggingface.co/Qwen/Qwen3-Embedding-8B in the MTEB learderboard. The encoder uses a 4096-dimensional embedding space with last-token pooling. We fine-tune the Qwen3-8B on the C-STS dataset with Pearson correlation loss computed over the cosine similarities. For the LoRA configuration, we use a rank $r=32$, a scaling factor $\alpha=32$, and a dropout rate of $0.05$. Adaptation is applied to the query, key, value, and output projection layers of the attention mechanism. Fine-tuning Qwen3-8B requires over four hours of wall-clock time on two NVIDIA A100 (80GB) GPUs. In contrast, CASE learns a supervised FFN on pre-computed embeddings and completes training in under one minute on a single NVIDIA RTX A6000 GPU (approximately 10 seconds on an A100). Moreover, LoRA fine-tuning of 8B-scale models requires substantial memory for model loading and activation storage, often necessitating at least 24 GB of GPU memory in practical settings, while CASE supervision requires negligible GPU memory usage (<1 GB).
[…]
@@
We propose CASE, a method for computing the semantic textual similarity between two sentences under a given condition. CASE encodes the condition under a sentence, and then subtracts the embedding of the condition in a post-processing step. Our experimental results show that LLM embeddings consistently outperform MLM embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to improve the performance in C-STS while projecting embeddings to lower-dimensional spaces for efficient similarity computations.
Acknowledgements
Danushka Bollegala holds concurrent appointments as a Professor at the University of Liverpool and as an Amazon Scholar at Amazon. This paper describes work performed at the University of Liverpool and is not associated with Amazon.